# Unbiased Regression-Adjusted Estimation of Average Treatment Effects in Randomized Controlled Trials

### v1

@@ 3 Regression Adjustment for Horvitz-Thompson Estimator
It is well known that, under simple random assignment, the classical Horvitz–Thompson (HT) estimator is unbiased for $\tau$ horvitz1952generalization. The variance of the HT estimator under simple random assignment is given by
@@
$\mu_{i}=\sqrt{\frac{1-p_{i}}{p_{i}}}\,y^{(1)}_{i}+\sqrt{\frac{p_{i}}{1-p_{i}}}\,y^{(0)}_{i}$. (2)
@@
$\breve{y}^{(1)}_{i}=y^{(1)}_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}\quad\text{and}\quad\breve{y}^{(0)}_{i}=y^{(0)}_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}.$ (3)
@@
Equation (4) implies that regression adjustment can reduce the variance of the HT estimator. In practice, however, only one potential outcome is observed for each unit, so $\boldsymbol{\mu}$} cannot be computed directly, and the minimization problem in (5) cannot be solved exactly. We address this challenge in LOORA-HT by replacing $\boldsymbol{\mu}$} with a proxy vector equal to $\boldsymbol{\mu}$} in expectation.
@@ 3.1 Leave-One-Out Regression-Adjusted Horvitz-Thompson
Algorithm˜1 describes LOORA-HT. Rather than using a single value for $\bm{b}$}, LOORA-HT applies a different vector of regression coefficients, $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$, for each unit $i$ in the sample. LOORA-HT constructs its adjustment vectors using a proxy vector $\widetilde{{\bm{y}}}$ that coincides with $\boldsymbol{\mu}$} in expectation,
@@
Each $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$ is obtained from a regression on the leave-one-out data, $(\widetilde{{\bm{y}}}_{-i},\widetilde{{\bm{X}}}_{-i})$. Leave-one-out coefficients can be particularly unstable for rows with high leverage scores (see Lemma˜B.1). To address this challenge, we use LOORA-HT ridge regression to mitigate the impact of high leverage scores. In Lemma˜2 below, we provide an upper bound on the ridge leverage scores as a function of the regularization parameter. For datasets with low leverage scores, the regularization parameter, $\lambda$, can be set to zero.
LOORA-HT is equal to
$\widehat{\tau}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}).$
@@ 3.2 The Variance of LOORA-HT
which characterizes how far the vector $\widetilde{{\bm{y}}}$ deviates from $\boldsymbol{\mu}$}. More precisely,
$\widetilde{{\bm{y}}}-{\boldsymbol{\mu}}=\frac{{\bm{z}}{\bm{t}}}{\boldsymbol{q}},$
@@
The LOORA-HT estimator (Algorithm˜1) is unbiased and its variance is given by the following formula:
@@
Examining (7) closely, we observe that its first term mirrors the regression-adjusted version of the variance of the classical HT estimator in (4), and quantifies the variance reduction achieved by regression adjustment. The factor $(1-\widetilde{h}_{\lambda ii})^{-2}$ arises from the removal of a single row in the leave-one-out regressions. More specifically, by Lemma˜B.1, the first term of (7) can be written as
@@
Note that when a leverage score $\widetilde{h}_{\lambda ii}$ equals one, the variance becomes unbounded. In such high-leverage cases, it is essential to employ ridge regression with a sufficiently large regularization parameter $\lambda$. The following lemma provides an upper bound on the ridge leverage scores.
@@
In our empirical analysis of within-subject experimental applications (Section˜6), we set $c=2$. It is well documented in the literature that units with high leverage scores can cause regression-adjusted estimators to yield unreliable estimates of the ATE (young2019channeling). Our LOORA estimators provide a systematic approach to mitigating this issue.
The second term of (7) arises from using the random vector $\widetilde{{\bm{y}}}$ in place of $\boldsymbol{\mu}$} in our regressions, since $\boldsymbol{\mu}$} is unobserved. As discussed earlier, ${\mathbb{E}}[\widetilde{{\bm{y}}}]={\boldsymbol{\mu}}$ and $\widetilde{{\bm{y}}}={\boldsymbol{\mu}}+\frac{{\bm{z}}{\bm{t}}}{\boldsymbol{q}}$. The second term of (7) thus captures the error introduced by estimating $\boldsymbol{\beta}_{\lambda}^{(-i)}$ with $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$, the solution to the regression based on $\widetilde{{\bm{y}}}$.
@@
In particular, for $\lambda=\left\|{\bm{X}}\right\|_{2,\infty}^{2}$, by Lemma˜2, the variance of LOORA-HT is bounded by
@@
Increasing the value of $\lambda$ decreases the second term of (7). It also reduces the magnitude of $(1-\widetilde{h}_{\lambda ii})^{-2}$, but at the same time increases the quantity
$\frac{1}{n^{2}}\sum_{i=1}^{n}(\widetilde{{\bm{x}}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}-\mu_{i})^{2}.$
Since the vector $\mu$ is not observed, one cannot search for the optimal value of $\lambda$.
We next study the asymptotic behavior of the LOORA-HT estimator. Under mild regularity and uniform boundedness assumptions, the variance of the LOORA-HT estimator with $\lambda=0$ converges as follows:
${\rm Var}\bigl(\sqrt{n}\,(\widehat{\tau}_{\textup{LHT}}-\tau)\bigr)\;\longrightarrow\;\lim_{n\to\infty}\frac{1}{n}\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}^{*}-{\boldsymbol{\mu}}\right\|_{2}^{2},$
where $\boldsymbol{\beta}^{*}$ is the oracle coefficient vector defined in (5). This limiting variance coincides with the asymptotic variance of the classical regression adjustment estimator with interaction terms (lin2013agnostic).
@@ 3.3 Asymptotic Normality of LOORA-HT
In this section, we establish the asymptotic normality of the LOORA-HT estimator. We adopt the standard asymptotic framework under the Neyman–Rubin potential outcomes model, in which the sequence of units, together with their covariates and potential outcomes, is viewed as a deterministic infinite population. Formally, let $({\bm{x}}_{i},y_{i}^{(1)},y_{i}^{(0)})_{i\in\mathbb{N}}$ denote this sequence, and consider the truncated sample $({\bm{x}}_{i},y_{i}^{(1)},y_{i}^{(0)})_{i\in[n]}$ as $n\to\infty$. The estimator is analyzed on this growing sequence.
We assume that the number of covariates is fixed and impose a set of regularity conditions ensuring that covariates and potential outcomes are uniformly bounded and well-behaved in the limit. Relative to lin2013agnostic, we introduce two additional restrictions: (i) the covariates have uniformly bounded norms, and (ii) the individual treatment effects are uniformly bounded. The first condition guarantees that the ridge leverage scores converge uniformly to zero, while the second ensures that the additional variance component—namely, the second term of (7)—vanishes asymptotically. Note that these assumptions are mild, as otherwise the covariates and individual treatment effects would diverge as $n$ increases.
  1. Assumption 1.
@@
    There exists a finite constant $L<\infty$ such that, for all $i\in\mathbb{N}$,
    $\left\|{\bm{x}}_{i}\right\|_{2}\leq L,\qquad\left|y_{i}^{(1)}-y_{i}^{(0)}\right|\leq L.$
    Moreover, for all $n\in\mathbb{N}$ and $j\in[k]$,
    $\frac{1}{n}\sum_{i=1}^{n}(y_{i}^{(1)})^{4}\leq L,\qquad\frac{1}{n}\sum_{i=1}^{n}(y_{i}^{(0)})^{4}\leq L,\qquad\frac{1}{n}\sum_{i=1}^{n}x_{ij}^{4}\leq L.$
  2. Assumption 2.
    Let $\bm{X}$} denote the $n\times k$ matrix whose $i$-th row is ${\bm{x}}_{i}$. Then $n^{-1}{\bm{X}}^{\top}{\bm{X}}$ converges to a bounded, positive definite matrix. Furthermore, the population means of $y_{i}^{(1)}$, $y_{i}^{(0)}$, $(y_{i}^{(1)})^{2}$, $(y_{i}^{(0)})^{2}$, $y_{i}^{(1)}y_{i}^{(0)}$, $y_{i}^{(1)}{\bm{x}}_{i}$, and $y_{i}^{(0)}{\bm{x}}_{i}$ converge to finite limits. For instance,
    $\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(1)}{\bm{x}}_{i}$
    exists and is finite.
  3. Assumption 3.
    For simple random assignment (used with LOORA-HT), let $p_{i}$ denote the treatment probability for unit $i$. We assume $m:=\inf_{i\in\mathbb{N}}\min\{p_{i},1-p_{i}\}$ is bounded away from zero. For complete random assignment (used with LOORA-DM), we assume that the treatment proportion $n_{T}/n$ converges to a limit $p_{T}$ satisfying $0<p_{T}<1$, which is bounded away from both 0 and $1$.
Recall that the outcome for unit $i$ is adjusted using the coefficient vector
@@
$\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}=\arg\min_{{\bm{b}}\in\mathbb{R}^{d}}\left\|\widetilde{{\bm{X}}}_{-i}{\bm{b}}-\widetilde{{\bm{y}}}_{-i}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$ (8)
Furthermore, define the population limit
@@
$\boldsymbol{\beta}^{*}_{\lambda}:=\lim_{n\to\infty}\arg\min_{{\bm{b}}\in\mathbb{R}^{d}}\left\|\widetilde{{\bm{X}}}{\bm{b}}-{\boldsymbol{\mu}}\right\|_{2}^{2}+\lambda\left\|{\bm{b}}\right\|_{2}^{2}.$ (9)
The next lemma establishes the existence of $\boldsymbol{\beta}^{*}_{\lambda}$ and the convergence in probability of the leave-one-out estimators $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ to this limit.
@@
Suppose the potential outcomes, covariates, and treatment assignment probabilities for the LOORA-HT estimator satisfy Assumptions˜1, 2 and 3. Let $\lambda\geq 0$ be fixed. Then the limit $\boldsymbol{\beta}^{*}_{\lambda}$ defined in (9) exists, and for all $i\in\mathbb{N}$,
$\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\;\xrightarrow{p}\;\boldsymbol{\beta}^{*}_{\lambda}.$
The previous lemma establishes that the leave-one-out coefficients $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$ converge in probability to their population limit $\boldsymbol{\beta}^{*}_{\lambda}$. This result ensures that, asymptotically, each unit’s adjusted outcome behaves as if it were constructed using the deterministic population parameter $\boldsymbol{\beta}^{*}_{\lambda}$ rather than the sample-dependent estimate $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}$. Consequently, the random variation introduced by finite-sample estimation of regression adjustments vanishes in large samples. Building on this observation, we can now characterize the asymptotic distribution of the LOORA-HT estimator itself.
@@
Under Assumptions˜1, 2 and 3, $\sqrt{n}({\widehat{\tau}}_{\textup{LHT}}-\tau)$ converges in distribution to a Gaussian random variable with mean 0 and variance
$\lim_{n\to\infty}\frac{1}{n}\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}^{*}-{\boldsymbol{\mu}}\right\|_{2}^{2}.$
Theorem 4 establishes the large-sample distribution of the LOORA-HT estimator and identifies its asymptotic variance in closed form. This result shows that, under mild regularity conditions, the regression adjustment embedded in LOORA-HT not only preserves consistency but also yields an asymptotically efficient estimator whose variance matches the limiting expression $\lim_{n\to\infty}n^{-1}\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}^{*}-{\boldsymbol{\mu}}\|_{2}^{2}$. In the next subsection, we build on this characterization to analyze the asymptotic efficiency of LOORA-HT and show that its limiting variance coincides with that of the classic regression-adjusted estimator with interaction terms.
3.4 Asymptotic Efficiency of LOORA-HT
We now examine the asymptotic efficiency of the LOORA-HT estimator. Theorem 4 established that $\widehat{\tau}_{\textup{LHT}}$ is asymptotically normal with a well-defined limiting variance. In this subsection, we show that under simple random assignment, the asymptotic variance of LOORA-HT coincides with that of the regression-adjusted estimator with interaction terms proposed by lin2013agnostic for complete random assignment. This result implies that LOORA-HT achieves the same asymptotic efficiency as Lin’s estimator, despite being constructed under a different assignment mechanism and without explicitly introducing interaction terms between treatment indicators and covariates.
The distinction between simple and complete random assignment is largely asymptotically negligible, but it matters for finite samples and for the structure of the adjustment. Lin’s specification augments the regression model with treatment–covariate interaction terms, effectively doubling the number of covariate coefficients to be estimated. While this ensures consistency and efficiency in large samples, it can also increase the dimensionality of the regression problem and lead to instability when the number of covariates is large relative to the sample size. By contrast, LOORA-HT achieves the same asymptotic efficiency through a leave-one-out regularized adjustment that avoids interactions and remains computationally stable even with high-dimensional covariates.
lin2013agnostic establishes the following asymptotic variance for the regression-adjusted estimator with interaction terms, denoted by $\widehat{\tau}_{\text{interact}}$:
$\lim_{n\to\infty}{\rm Var}\!\big(\sqrt{n}(\widehat{\tau}_{\text{interact}}-\tau)\big)=\frac{1-p_{T}}{p_{T}}\lim_{n\to\infty}\sigma_{T}^{2}+\frac{p_{T}}{1-p_{T}}\lim_{n\to\infty}\sigma_{C}^{2}+2\lim_{n\to\infty}\sigma_{TC},$ (10)
where $p_{T}$ is the limiting treatment proportion and
@@ 3.4 Asymptotic Efficiency of LOORA-HT
$\sigma_{T}^{2}$ $:=\frac{1}{n}\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{T}^{*}-({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})\right\|_{2}^{2},$
$\sigma_{C}^{2}$ $:=\frac{1}{n}\left\|({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{C}^{*}-({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\right\|_{2}^{2},$
$\sigma_{TC}$ $:=\frac{1}{n}\Big[\big(({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{T}^{*}-({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})\big)^{\top}\big(({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{C}^{*}-({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\big)\Big],$
and
$\boldsymbol{\beta}_{T}^{*}$ $:=\lim_{n\to\infty}\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}}\left\|({\bm{X}}-\overline{{\bm{X}}}){\bm{b}}-({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})\right\|_{2}^{2},$
$\boldsymbol{\beta}_{C}^{*}$ $:=\lim_{n\to\infty}\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}}\left\|({\bm{X}}-\overline{{\bm{X}}}){\bm{b}}-({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\right\|_{2}^{2}.$
We now express the right-hand side of (10) in a form that facilitates comparison with the asymptotic variance of LOORA-HT. Straightforward algebra yields
$\frac{1-p_{T}}{p_{T}}\sigma_{T}^{2}+\frac{p_{T}}{1-p_{T}}\sigma_{C}^{2}+2\sigma_{TC}$
$=\frac{\big\|(1-p_{T})\big(({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{T}^{*}-({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})\big)+p_{T}\big(({\bm{X}}-\overline{{\bm{X}}})\boldsymbol{\beta}_{C}^{*}-({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\big)\big\|_{2}^{2}}{n\,p_{T}(1-p_{T})}$
$=\frac{\big\|({\bm{X}}-\overline{{\bm{X}}})\!\big((1-p_{T})\boldsymbol{\beta}_{T}^{*}+p_{T}\boldsymbol{\beta}_{C}^{*}\big)-\big((1-p_{T})({\bm{y}}^{(1)}-\overline{{\bm{y}}}^{(1)})+p_{T}({\bm{y}}^{(0)}-\overline{{\bm{y}}}^{(0)})\big)\big\|_{2}^{2}}{n\,p_{T}(1-p_{T})}$
$=\frac{1}{n}\big\|({\bm{X}}-\overline{{\bm{X}}})\big(\sqrt{\tfrac{1-p_{T}}{p_{T}}}\boldsymbol{\beta}_{T}^{*}+\sqrt{\tfrac{p_{T}}{1-p_{T}}}\boldsymbol{\beta}_{C}^{*}\big)-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\big\|_{2}^{2}.$
Moreover, one can verify that
$\sqrt{\tfrac{1-p_{T}}{p_{T}}}\boldsymbol{\beta}_{T}^{*}+\sqrt{\tfrac{p_{T}}{1-p_{T}}}\boldsymbol{\beta}_{C}^{*}=\lim_{n\to\infty}\mathop{\mathrm{argmin}\,}\limits_{{\bm{b}}}\left\|({\bm{X}}-\overline{{\bm{X}}}){\bm{b}}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\right\|_{2}^{2}.$
To connect this expression with LOORA-HT, consider the case where $p_{i}=p_{T}$ for all $i\in\mathbb{N}$. In this setting,
$\widetilde{{\bm{X}}}(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}})^{-1}\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}}={\bm{X}}({\bm{X}}^{\top}{\bm{X}})^{-1}{\bm{X}}^{\top}{\boldsymbol{\mu}}.$ (11)
To compare the asymptotic variance of LOORA-HT with that of $\widehat{\tau}_{\text{interact}}$, we set $\lambda=0$. Using standard calculations (see, e.g., Appendix A of ghadiri2023finite), if we apply Algorithm˜1 with the matrix $\big[\,\widetilde{{\bm{X}}}-\overline{\widetilde{{\bm{X}}}}\,\big|\,\textbf{1}\,\big],$ that is, centering the covariates and including an intercept, then the variance of $\sqrt{n}(\widehat{\tau}_{\textup{LHT}}-\tau)$ converges to
$\lim_{n\to\infty}\frac{1}{n}\min_{{\bm{b}}}\left\|(\widetilde{{\bm{X}}}-\overline{\widetilde{{\bm{X}}}}){\bm{b}}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\right\|_{2}^{2}=\lim_{n\to\infty}\frac{1}{n}\min_{{\bm{b}}}\left\|({\bm{X}}-\overline{{\bm{X}}}){\bm{b}}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\right\|_{2}^{2},$
where the equality follows from (11). Therefore, the asymptotic variance of LOORA-HT is equal to that of $\widehat{\tau}_{\text{interact}}$.
This equivalence demonstrates that LOORA-HT attains the same asymptotic efficiency as the regression-adjusted estimator with interaction terms of lin2013agnostic, even though it is constructed under simple random assignment and does not rely on explicitly modeling treatment–covariate interactions. In the next subsection, we develop valid and asymptotically correct confidence intervals for LOORA-HT.

### v2

@@ 3 Regression Adjustment for Horvitz-Thompson Estimator
It is well known that, under simple random assignment, the classical Horvitz–Thompson (HT) estimator is unbiased for $\tau$. The variance of the HT estimator under simple random assignment is given by
@@
$\mu_{i}=\sqrt{\frac{1-p_{i}}{p_{i}}}\,y^{(1)}_{i}+\sqrt{\frac{p_{i}}{1-p_{i}}}\,y^{(0)}_{i}$.
@@
$\breve{y}^{(1)}_{i}=y^{(1)}_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}\quad\text{and}\quad\breve{y}^{(0)}_{i}=y^{(0)}_{i}-{\bm{x}}_{i}^{\top}{\bm{b}}.$
@@
Equation (2) implies that regression adjustment can reduce the variance of the HT estimator. In practice, however, only one potential outcome is observed for each unit, so $\boldsymbol{\mu}$} cannot be computed directly, and the minimization problem in (3) cannot be solved exactly. We address this challenge in LOORA-HT by replacing $\boldsymbol{\mu}$} with a proxy vector that is equal to $\boldsymbol{\mu}$} in expectation.
@@ 3.1 Leave-One-Out Regression-Adjusted Horvitz-Thompson
Algorithm˜1 describes LOORA-HT. Rather than using a single value for $\bm{b}$}, LOORA-HT applies a different vector of regression coefficients, $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$, to each unit $i$ in the sample. LOORA-HT constructs its adjustment vectors using a proxy vector $\widetilde{{\bm{y}}}$ that coincides with $\boldsymbol{\mu}$} in expectation,
@@
LOORA-HT is defined as
$\widehat{\tau}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}),$
where $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$ is obtained from a regression on the leave-one-out data, $(\widetilde{{\bm{y}}}_{-i},\widetilde{{\bm{X}}}_{-i})$. In general, leave-one-out coefficients can be unstable, particularly for rows with high leverage scores (see Lemma˜6). To address this challenge, we use ridge regression to mitigate the impact of high leverage scores. In Lemma˜1 below, we provide an upper bound on the ridge leverage scores as a function of the ridge regularization parameter.
@@ 3.2 The Variance of LOORA-HT
which characterizes how far $\widetilde{{\bm{y}}}$ deviates from $\boldsymbol{\mu}$},
$\widetilde{{\bm{y}}}-{\boldsymbol{\mu}}=\frac{{\bm{z}}{\bm{t}}}{\boldsymbol{q}}.$
@@
Under simple random assignment, LOORA-HT (Algorithm˜1) is exactly unbiased, with variance equal to
@@
The first term in (5) mirrors the variance of the infeasible regression-adjusted version HT estimator in (2), and quantifies the variance reduction achieved by regression adjustment. The factor $(1-\widetilde{h}_{\lambda ii})^{-2}$ arises from the removal of a single row in the leave-one-out regressions. More specifically, by Lemma˜6, the first term of (5) can be written as
@@
When a leverage score $\widetilde{h}_{\lambda ii}$ approaches one, the variance becomes unbounded. In such high-leverage cases, it is essential to employ ridge regression with a sufficiently large regularization parameter $\lambda$. The following lemma provides an upper bound on the ridge leverage scores.
@@
High-leverage observations distort regression-adjusted estimates. LOORA applies ridge regularization, which stabilizes the adjustment and delivers more reliable estimates.
The second term of (5) arises from using the random vector $\widetilde{{\bm{y}}}$, with ${\mathbb{E}}[\widetilde{{\bm{y}}}]={\boldsymbol{\mu}}$, in place of $\boldsymbol{\mu}$} in the regression adjustment. In other words, it captures the error introduced by estimating $\boldsymbol{\beta}_{\lambda}^{(-i)}$ with $\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}$.
@@
In particular, for $\lambda=\left\|\widetilde{{\bm{X}}}\right\|_{2,\infty}^{2}$, by Lemma˜1, the variance of LOORA-HT is bounded by
@@
We next study the asymptotic behavior of the LOORA-HT estimator.
@@ 3.3 Asymptotic Normality of LOORA-HT
This section derives a large-sample approximation to the distribution of LOORA-HT under the following assumptions.
  1. Assumption 1
@@
    (uniform boundedness). There exists a finite constant $L<\infty$ such that, for all $i\in\mathbb{N}$, $\left\|{\bm{x}}_{i}\right\|_{2}\leq L$, $|y_{i}^{(1)}|\leq L$, and $|y_{i}^{(0)}|\leq L$.
  2. Assumption 2
    (positivity). There exists a constant $m\in(0,0.5]$ such that for simple random assignment $m<p_{i}<1-m$, for all $i\in\mathbb{N}$; and for complete random assignment $m<n_{T}/n<1-m$ for all $n\in\mathbb{N}$.
  3. Assumption 3
    (positive definiteness). The smallest eigenvalue of $n^{-1}{\bm{X}}^{\top}{\bm{X}}$ is bounded away from zero uniformly in $n$.
Assumptions 1 and 3 provide bounds on potential outcomes and covariates, but allow the means of $(y_{i}^{(1)},y_{i}^{(0)},x_{i})$ to fluctuate, without necessarily converging to a limit as $n$ increases. Assumption 2 requires that each experimental unit be assigned to treatment and control with strictly positive probabilities. It rules out degenerate experimental designs in which causal effects cannot be identified.
Let
$\boldsymbol{\beta}_{\lambda}=(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}{\boldsymbol{\mu}}$
be the true solution of the ridge regression problem on $\boldsymbol{\mu}$}. The next theorem establishes the large-sample distribution of the LOORA-HT estimator and identifies its asymptotic variance in closed form.
Theorem 2.
Suppose Assumptions˜1, 2 and 3 hold. Let $\lambda\geq 0$ be fixed. Then, if $\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\|_{2}\to\infty$, under simple random assignment,
@@ 3. Assumption 3
$\frac{\sqrt{n}({\widehat{\tau}}_{\textup{LHT}}-\tau)}{\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}/\sqrt{n}}$
converges in distribution to a standard normal random variable.
In the next subsection, we build on Theorem 2 to analyze the asymptotic efficiency of LOORA-HT.
3.4 Asymptotic Efficiency of LOORA-HT
We now examine the asymptotic efficiency of the LOORA-HT estimator. Theorem 2 established the asymptotic normality of LOORA-HT. In this subsection, we show that under simple random assignment, LOORA-HT is asymptotically efficient among the class of Horvitz-Thompson linearly adjusted estimators, defined as estimators of the form,
${\widehat{\tau}}_{\textup{HLA}}=\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\bigl(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\bigr),$
where there exists a bounded deterministic sequence of vectors $\boldsymbol{\gamma}_{1},\boldsymbol{\gamma}_{2},\ldots$ and a sequence of vectors $\widehat{\boldsymbol{\gamma}}_{1},\widehat{\boldsymbol{\gamma}}_{2},\ldots$ such that
$\max_{i\in[n]}\left\|\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}-\widehat{\boldsymbol{\gamma}}_{n}\right\|_{2}=o_{p}(n^{-1/2})\quad\text{and}\quad\left\|\boldsymbol{\gamma}_{n}-\widehat{\boldsymbol{\gamma}}_{n}\right\|_{2}\xrightarrow{p}0.$ (6)
The class of HT linearly adjusted estimators is larger than the class of estimators usually considered for the purpose of determining efficiency (e.g., see Cytrynbaum (2024)), because it allows for unit-dependent adjustment vectors, $\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}$.
LOORA-HT is contained in the class of HT linearly adjusted estimators since
@@ 3.4 Asymptotic Efficiency of LOORA-HT
$\max_{i\in[n]}\left\|\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}-\widehat{\boldsymbol{\beta}}_{\lambda}\right\|_{2}=O(n^{-1})\quad\text{and}\quad\left\|\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}_{\lambda}\right\|_{2}=O_{p}(n^{-1/2}),$ (7)
where
$\widehat{\boldsymbol{\beta}}_{\lambda}=(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{y}}}$,
as shown in the proof of Theorem˜2. The next result describes the asymptotic behavior of HT linearly adjusted estimators.
@@
Suppose Assumptions˜1, 2 and 3 hold and $\left\|\widetilde{{\bm{X}}}\boldsymbol{\gamma}_{n}-{\boldsymbol{\mu}}\right\|_{2}=\Omega(\sqrt{n})$. Then under simple random assignment, ${\widehat{\tau}}_{\textup{HLA}}$ is a consistent estimator and
$\frac{\sqrt{n}({\widehat{\tau}}_{\textup{HLA}}-\tau)}{\frac{1}{\sqrt{n}}\left\|\widetilde{{\bm{X}}}\boldsymbol{\gamma}_{n}-{\boldsymbol{\mu}}\right\|_{2}}\xrightarrow{d}\mathcal{N}(0,1).$
Using this result, we can establish the asymptotic efficiency of LOORA-HT in the case where $\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{0}-{\boldsymbol{\mu}}\right\|_{2}=\Omega(\sqrt{n})$.
@@
Under the assumptions of Lemma˜2 and simple randomization, LOORA-HT with $\lambda=0$ is an asymptotically efficient estimator in the class of HT linearly adjusted estimators.
For the case of $\lambda=0$ and provided that the regression adjustment includes an intercept, by Lemma˜5, the variance of LOORA-HT reduces to
@@
$\min_{\boldsymbol{\beta}\in{\mathbb{R}}^{k}}\frac{1}{n^{2}}\left\|(\widetilde{{\bm{X}}}-\overline{\widetilde{{\bm{X}}}})\boldsymbol{\beta}-({\boldsymbol{\mu}}-\overline{{\boldsymbol{\mu}}})\right\|_{2}^{2},$
This expression is similar to the variance of the estimator of Lin (2013).
In the next subsection, we develop valid and asymptotically correct confidence intervals for LOORA-HT.